# Structure and Control of Biology-inspired Networks

### v1

@@ 1 Introduction
Literature Review
A brief background survey of bio-inspired and neuro-inspired control theory could divide current research into two broad areas: the micro case and the macro case. The micro case considers systems with a large number of simple control inputs that are coordinated to achieve prescribed actions. In this context, Beyeler et al. (2019) viewed such simple inputs as neuronal responses, which can be understood as an emergent property of nonnegative sparse coding (NSC). NSC might be utilized by sensory cortical regions to efficiently encode representations of external stimuli, while basal ganglia may employ NSC for the coordination of motor behaviors. Building upon this notion of simple neuronal inputs, Baillieul (2019) proposed neuromimetic control models that utilize such inputs collectively, associating them with system-wide objectives. Then, Sun and Baillieul (2022) aimed to elucidate the dynamic characteristics of control systems actuated by the aggregate effects of inputs — simulating neural states of spiking or resting. Following this line of investigation, Sun and Baillieul (2024) explored neural heuristic quantization systems to learn quantized motions and develop resilience to channel dropouts. It was demonstrated that by employing algorithms akin to Oja’s rule by Oja (1982), these neuromimetic models could emulate classical control designs produced by a large number of analog inputs. The macro case, on the other hand, investigates biologically plausible implementations coupled with different learning rules, such as the Hebbian rules introduced by Galtier et al. (2012), which were shown to be capable of learning geometric features of cortical organization., and Oja’s rules explored by Pehlevan et al. (2015), which focused on unsupervised learning to solve a linear dimensionality reduction problem. An extension work by Pehlevan and Chklovskii (2019) demonstrated that neural networks equipped with both Hebbian and anti-Hebbian learning rules can perform a broad range of unsupervised learning tasks. Meanwhile, Krotov and Hopfield (2019) have proposed a novel biologically plausible learning rule, demonstrating its efficacy on the MNIST dataset. Recently, Centorrino et al. (2024) provided an overview of these models and combined the Hopfield Neural Network and Firing Rate Neural Network with two different Hebbian learning rules, analyzing the contractivity of each coupled model by leveraging non-Euclidean contraction arguments. Indeed, when analyzing dynamic models, a crucial problem that arises is ensuring the stability and controllability of the system; see Gu et al. (2015); Menara et al. (2018a). However, the criterion to prove whether a system is controllable can be stringent at times, as the coefficient matrices are frequently unknown. To address this challenge, the concept of structural controllability, first proposed by Lin (1974), has been employed in analyzing the dynamics of linear and linearized brain system models (Menara et al. (2018b); Lynn and Bassett (2019); Kim and Bassett (2020)). It relaxes the requirement of precise knowledge of the system parameters and instead focuses on the underlying structure of the system, determined by the nonzero entries in the system matrices. Beyond structural controllability, the key feature of any brain network model that is capable of simple decision tasks like classifying data is nonliearity (Ruiz-Balet and Zuazua (2023)), and the introducing nonlinearity to the control theory of network systems is a key contribution of the present paper.
@@
In the context of the above literature, the following sections propose network-based models aimed at reproducing biological capabilities such as learning and predictive cognition. The frameworks focus both on individual neurons and on interconnected brain regions. This paper offers four key contributions: (i) we propose a coupled pre- and post-synaptic model undergoing Hebbian learning and anti-Hebbian rules, capturing state-dependent synaptic plasticity in both excitatory and inhibitory connections. The updating rules for the weights of synapses use only local information based on the neuron connections; (ii) we leverage graph theory and control theory to analyze this model and derive conditions ensuring reflection of neuronal characteristics, such as stability and structural controllability; (iii) we give sufficient conditions based on biologically meaningful quantities, such as synaptic decay rate and topological connectivity that preserve neural properties; (iv) finally, we validate theoretical results via simulations that include the dynamics of a model monkey brain parcellation.
@@ 2.1 Graph theory and Structure Patterns of Linear Systems
In the present paper, we shall be interested in a special class of graphs known as cactus graphs. These have proven to be useful in control and network theory in a number of ways—particularly as related to dissipativity and stability. We refer to Arcak (2011) for a number of recent references. The work that follows constructively enlarges much earlier work that was initiated by Lin (1974), and to develop the ideas, we adopt some of Lin’s terminology.
@@
(Bud) A bud is a subgraph containing a cycle and an additional node that is connected by a single edge to a node in the cycle. We call this connecting edge the distinguished edge and the additional node dangling node.
A number of authors, including Lin (1974); Rech and Perret (1990); Mayeda (1981) have studied buds with directed distinguished edge connections. In these papers, a distinction is made between input bud and output bud, where the edge is directed from the cycle or the edge is directed toward the cycle. In this paper, we focus on the bidirectional bud, which refers to both input and output perspectives, with a bidirectional distinguished edge between the cycle and dangling node. Keeping with the botanical terminology, Lin referred to a simple path to which buds are attached as a stem. For each bud, we refer to the endpoint of the distinguished edge that is not on the cycle as a joint, and in Lin’s definition of a cactus graph, we refer to the nodes located on the stem that connect the bud to it as the joints.
@@
(Cactus) Given a stem $S$ and buds $B_{1},B_{2},\dots,B_{n}$, the structure $S\cup B_{1}\cup\dots\cup B_{n}$ is called a cactus if for all $i\in\{1,2,\dots n\}$, $\{S\cup B_{1}\cup\dots\cup B_{i-1}\}\cap B_{i}=v_{b_{i}}$, where $v_{b_{i}}$ is the dangling node of bud $B_{i}$.
The cactus structure described herein aligns with the definition provided in Lin (1974) and Menara et al. (2018a). It is important to note that this conceptualization differs slightly from the definition of a cactus tree in graph theory, where any two simple cycles share at most one common vertex. We say that a graph $\mathcal{G}_{a}=(\mathcal{V}_{a},\mathcal{E}_{a})$ is spanned by a graph $\mathcal{G}_{b}=(\mathcal{V}_{b},\mathcal{E}_{b})$ when $\mathcal{V}_{a}=\mathcal{V}_{b}$ and $\mathcal{E}_{b}\subseteq\mathcal{E}_{a}$. $\mathcal{G}_{a}$ and $\mathcal{G}_{b}$ can be connected or disconnected.
@@ 2.2 System Graphs and Structure Patterns
3 Modelling of Neuromimetic Dynamic Networks
@@
where $t\in(\tau p,\tau(p+1))$, $p=1,2,\cdots$ and $\tau$ is a time constant denoting the time interval of updating the edge weight. $i,j\in\mathcal{V}$. In (1a), $b_{i}=0$ means there is no external input injected to node $i$. The entry $a_{ij}$ represents the connection weight of the pre-synaptic neuron $j$ to the post-synaptic neuron $i$.
The activation function $\gamma_{\theta}$ is designed to emulate neuronal spiking behavior. It is defined as:
@@
where $\theta$ represents the threshold for neuronal activation. It can also be expressed using the standard $ReLU$ function, $\sigma(x)=\max(x,0)$ as $\gamma_{\theta}(x)=-\sigma(-x-\theta)+\sigma(x-\theta)+sign(\sigma(x-\theta))\theta+sign(-\sigma(-x-\theta))\theta$. While various activation functions exist, this serves as an illustrative example. If $x$ represents the membrane potential of the postsynaptic neuron, the function $\gamma_{\theta}(x)$ models the instantaneous increase in postsynaptic membrane potential caused by the presynaptic release of vesicles, under the condition that the threshold potential $+\theta/-\theta$ is reached. This models the scenario where presynaptic vesicle release results in a discontinuous jump in postsynaptic potential upon crossing the threshold $\theta/-\theta$, which corresponds to the integrate-and-fire model in Burkitt (2006).
@@ 4.1 Boundedness and Stability
In models of neurodynamics, a commonly used activation function is the sigmoid, which satisfies the boundedness assumption. In this paper, we use the range of the sigmoid function to be $[-1,1]$ such that $\sup{\phi}=1$ and $\phi(0)=0$. From the dynamic equations (2), we define $\mathcal{X}=\{\bm{x}\in\mathbb{R}^{n}:|x_{i}|\leq x_{max}\}$, where $x_{max}=\frac{||B\bm{u}||_{\infty}}{c_{n}-d^{in}\max\{\overline{a}^{+},|\underline{a}^{-}|\}}$ denotes the maximum membrane potential.
@@
Proof.
Denote the $p-$th time slot by $(p\tau,(p+1)\tau],p=0,1,2,\dots$, and let $t_{p}\in(p\tau,(p+1)\tau]$ be an arbitrary time instant in the time slot. $A_{p}$ is the corresponding weight matrix during this period, which is constant according to (2b). Therefore, $A(t)$ is a piecewise constant matrix with respect to $t$ and with all diagonal entries equal to 0. $C_{n}$ is a diagonal matrix with all diagonal entries satisfying the assumptions of Lemma 2. Consider any time slot $p$, and let $\tilde{H}_{p}=-C_{n}+A_{p}$. With $\bm{u}=\bm{0}$, the system (2a) is piecewise affine with respect to $\bm{x}$, with different definitions across regions where the affine function $\Gamma(A_{p}\bm{x})$ takes on different values determined by the thresholds ±$\pm\theta$. Enumerate the distinct homogenous terms of the affine pieces of $-C_{n}\bm{x}+\Gamma(A_{p}\bm{x})$ as $H_{k}\bm{x}$, $k=1,\dots,K_{p}$. Then for all $H_{k}$’s in the list we have $|H_{k}|\preceq|\tilde{H}_{p}|$, where the notation ||$\cdot$|| ⋅ | means that absolute values of each entry in $H_{k}$ is less than or equal to the absolute value of the corresponding entry in $\tilde{H}_{p}$. Only when all entries of $A_{p}\bm{x}$ are larger than $\theta$ or smaller than $-\theta$ does equality hold.
Since $\sum_{j=1,j\neq i}^{N}|a_{ij}|\leq d^{in}\max\{\overline{a}^{+},|\underline{a}^{-}|\}$, when conditions in lemma 2 are satisfied, for each row $i$ of $-\tilde{H}_{p}$, $-\tilde{H}_{p}(i,i)=c_{n}>d^{in}\max\{\overline{a}^{+},|\underline{a}^{-}|\}\geq\sum_{j=1,j\neq i}^{N}|a_{ij}|=\sum_{j=1,j\neq i}^{N}|-\tilde{H}_{p}(i,j)|$. Therefore, $-\tilde{H}_{p}$ is a Hermitian strict diagonally dominant matrix. Since each $|H_{k}|\preceq|\tilde{H}_{p}|$ and $H_{k}$ has the same diagonal entries as $\tilde{H}_{p}$, $H_{k}$ is also a Hermitian strict diagonally dominant matrix. By the Gershgorin circle theorem, we obtain $-H_{k}$ is positive definite and $H_{k}$ is Hurwitz for all $k$. Thus, there is a sequence of time steps $\epsilon_{k}$ such that the closed-form solution of the state in time slot $(p\tau,(p+1)\tau]$ can be be written explicitly for an appropriate ordering of indices $k$ as
$\begin{split}\bm{x}(t_{p+1})=e^{H_{K}\epsilon_{K}}\ \cdots\ e^{H_{1}\epsilon_{1}}\ \bm{x}(t_{p}).\\ \end{split}$ (4)
Since the $H_{k}$’s, are Hurwitz for all $k$, the solutions over time slots $[p\tau,(p+1)\tau)$ evolve toward 00 as $p\to\infty$ which implies (2a) is asymptotically stable. ∎
@@ 4.2 Structural Controllability and Observability
4.3 Minimal Control Energy
The determination of energy cost is of interest because the neural systems governing motor control in biological organisms are intrinsically inclined to identify and select movement strategies that minimize energy consumption, while still allowing for the successful execution of intended motions and actions. Denote the optimal control input that steers the system described by the model equation (2a) from an initial state $\bm{x_{0}}$ to a target state $\bm{x_{f}}$ by $\bm{u}$. The corresponding cost functional to be minimized is
[…]
Lemma 5.
Consider the system (2) with piecewise constant system matrix $A(t)$ and a generalized cactus structure with $K$ control input nodes. Suppose further that a path segmentation $\bm{x^{i}}$ is given and that a steering law $\bm{u}(t)$ has been chosen to steer $\bm{\dot{x}}=A(t)\bm{x}+Bu$ from $\bm{x_{p}}$ to $\bm{x_{p+1}}$ so as to minimize (6) over one time interval $t\in[p\tau,(p+1)\tau)$. If additional nodes have external inputs, the minimum energy cost $\tilde{\eta}_{p}$ corresponding to the optimal input $\bm{\hat{u}}$ with the additional nodes applied to the same path segmentaion will be less than or equal to the cost $\eta_{p}$ associated with the original optimal control input $\bm{u}$ on each segment.
During each of these time intervals, the system has a fixed weight matrix. If an additional node in the system is provided with an external input, according to Baillieul (2019)[Theorem 2], the optimal energy cost $\tilde{\eta}_{p}$ denoting the optimal energy for the $p$-th $\tau$-interval will be less than or equal to the original energy cost $\eta_{p}$, i.e., $\tilde{\eta}_{p}\leq\eta_{p}$.
Theorem 6.
If there is an optimal path segmentation $\bm{x^{i}}$ such that the steering law is optimal on each segment (according to criterion (6)) and then if additional input channels come online, the optimal cost (6) of the system with the additional inputs is less than or equal to the original cost.
Proof.
The generalized cactus structure guarantees the structural controllability of the system. According to (2), the weight matrix in the system model is updated every $\tau$ time units and remains constant within each $\tau$-interval. As a result, the optimal input $\bm{u}$ and the corresponding optimal trajectory $\mathcal{T}$ in the whole interval can be divided into several parts, $\bm{u}_{p}$ and $\mathcal{T}_{p}$, where $p=0,1,2,\cdots$. Furthermore, it is guaranteed that at the end of each $\tau$-interval, both the original and the new systems will reach the same state when applying the optimal input $\bm{u}_{p}$ and $\bm{\tilde{u}}_{p}$ respectively. Consider the entirety of the time interval, the minimum energy cost of the system augmented with an additional input node $\hat{\eta}$ should be less than or equal to the aggregate of the optimal energy expenditures over each constituent $\tau$-interval, i.e., $\hat{\eta}\leq\sum_{p}\tilde{\eta}_{p}$. Because the global optimal solution over the entire time span should not exceed the sum of locally optimal solutions summed over the small time segments. Meanwhile, from Lemma 5, we have $\tilde{\eta}_{p}\leq\eta_{p}$ at each $p$-th interval. Therefore, $\hat{\eta}\leq\sum_{p}\tilde{\eta}_{p}\leq\sum_{p}\eta_{p}=\eta$. This proves the theorem. ∎
This result can be intuitively understood, as the additional control inputs provide more flexibility in optimizing the system’s behavior, potentially leading to a lower energy cost compared to the case with less control input nodes.
@@ 4.4 Resilience and Cascade
It is obtained graphically. From Theorem 4, we obtain that as long as the graph preserves the generalized cactus structure during dynamic evolutions, the system is structurally controllable and observable relative to the same input nodes. Therefore, even if there are neurons and connections that are intermittently available, the system is always structurally controllable and observable. Meanwhile, even if there are edges and nodes unavailable, conditions in lemma 2 still hold. Thus, the state matrix in (2a) is still a Hermitian diagonally dominant matrix and thus, the system remains stable. ∎

### v2

@@ 1 Introduction
Literature Review A brief background survey of bio-inspired and neuro-inspired control theory could divide current research into two broad areas: the micro case and the macro case. The micro case considers systems with a large number of simple control inputs that are coordinated to achieve prescribed actions. In this context, Beyeler et al. (2019) viewed such simple inputs as neuronal responses, which can be understood as an emergent property of nonnegative sparse coding (NSC). Building upon this notion of simple neuronal inputs, Baillieul (2019) and Sun and Baillieul (2022) proposed neuromimetic control models that utilize such inputs collectively, associating them with system-wide objectives. Following this investigation, Sun and Baillieul (2024) explored neural heuristic quantization and showed that quantized control input sequences acting in parallel through large numbers of input channels could achieve control objective while being resilient to channel dropouts. This line of research primarily focuses on how multiple neurons collaborate to accomplish tasks, closely resembling real microscopic biological models. However, it has certain limitations. For instance, when constructing control systems, directly integrating biologically plausible principles into system parameters proves challenging, often requiring the reconstruction of algorithms as a supplement for simultaneous application. Furthermore, while this approach is effective for tasks such as emulation, it faces difficulties in addressing a broader range of real-world challenges, including complex problems like data classification, path planning, and others.
The macro case, as treated in this paper, investigates biologically plausible implementations coupled with different learning rules, such as the Hebbian rules introduced by Galtier et al. (2012), which were shown to be capable of learning geometric features of cortical organization by extracting the geometry of input spaces using slow Hebbian learning in Hopfield networks, and Oja’s rules explored by Pehlevan et al. (2015), which proposes a single-layer neuronal network that performs online symmetric non-negative matrix factorization (SNMF) of the similarity matrix of streamed data. The algorithm is derived from a principled cost function and can be implemented by a biologically plausible network with local learning rules. Meanwhile, Krotov and Hopfield (2019) have proposed a novel biologically plausible learning rule that utilizes global inhibition in the hidden layer to extract key features, demonstrating its efficacy on the MNIST dataset. While all of the aforementioned papers are grounded in Hebbian learning and recurrent dynamics, each adopts a specific mathematical framework, such as SNMF, rather than providing a more comprehensive biological description. This methodological focus introduces certain limitations, as backpropagation algorithms may be replaced in initial stages, yet the scope of tasks that can be effectively addressed remains restricted. In some cases, this approach necessitates the addition of another layer trained with labeled data using conventional methods, such as stochastic gradient descent, as discussed by Krotov and Hopfield (2019).Furthermore, these studies generally lack an investigation into critical properties such as the controllability and stability of the system. Recently, a similar work with this paper by Centorrino et al. (2024) provided an overview of the biologically inspired models and combined the Hopfield Neural Network and Firing Rate Neural Network with two different Hebbian learning rules, analyzing the contractivity of each coupled model by leveraging non-Euclidean contraction arguments. While it primarily focuses on theoretical analysis without presenting practical application scenarios, our work extends beyond theoretical exploration. By leveraging graph theory and control theory, we analyze the system’s various biologically plausible properties. Additionally, we demonstrate the applicability of our approach through its implementation on a macaque visual neural network from Kandel et al. (2000) and its use in data classification tasks.
@@
In the context of the above literature, this paper offers four key contributions: (i) we propose a coupled pre- and post-synaptic model undergoing Hebbian learning and anti-Hebbian rules, capturing state-dependent synaptic plasticity in both excitatory and inhibitory connections. The updating rules for the weights of synapses use only local information based on the neuron connections; (ii) we leverage graph theory and control theory to analyze this model and derive conditions ensuring reflection of neuronal characteristics, such as stability, structural controllability and resilience; (iii) we give sufficient conditions based on biologically meaningful quantities, such as synaptic decay rate and topological connectivity that preserve neural properties; (iv) finally, we validate theoretical results via simulations that include the dynamics of a model monkey brain parcellation.
@@ 2.1 Graph theory and Structure Patterns of Linear Systems
In the present paper, we shall be interested in a special class of graphs known as cactus graphs. We refer to Arcak (2011) for a review of recent literature, which defines cactus graphs as structures where any two simple cycles share at most one common vertex, as illustrated in Fig. 1. However, it is important to note that such a structure alone does not constitute a sufficient condition for graph controllability. For instance, Fig. 1(a) depicts a graph that is not controllable, whereas Fig. 1(b) represents a controllable graph. Consequently, our subsequent work builds upon and significantly expands the research initiated by Lin (1974). To further develop these concepts, we adopt Lin’s definition of cactus graphs as a foundational framework for our analysis.
Figure 1: The figure illustrates two general cactus structures as described by Arcak (2011), with the structure on the right also exemplifying Lin’s concept of a (spanned) cactus graph.
@@
(Bud) A bud is a subgraph containing a cycle and an additional node that is connected by a single edge to a node in the cycle. We call this connecting edge the distinguished edge and the additional node joint node.
A number of authors, including Lin (1974); Rech and Perret (1990); Mayeda (1981) have studied buds with directed distinguished edge connections. In these papers, a distinction is made between input bud and output bud, where the edge is directed from the cycle or the edge is directed toward the cycle. In this paper, we focus on the bidirectional bud, which refers to both input and output perspectives, with a bidirectional distinguished edge between the cycle and joint node. Keeping with the botanical terminology, Lin referred to a simple path to which buds are attached as a stem.
@@
(Cactus) Given a stem $S$ and buds $B_{1},B_{2},\dots,B_{n}$, the structure $S\cup B_{1}\cup\dots\cup B_{n}$ is called a cactus if for all $i\in\{1,2,\dots n\}$, $\{S\cup B_{1}\cup\dots\cup B_{i-1}\}\cap B_{i}=v_{b_{i}}$, where $v_{b_{i}}$ is the joint node of bud $B_{i}$.
The cactus structure described herein aligns with the definition provided in Lin (1974) and Menara et al. (2018). It is important to note that this conceptualization differs slightly from the definition of a cactus tree in graph theory, where any two simple cycles share at most one common vertex. We say that a graph $\mathcal{G}_{a}=(\mathcal{V}_{a},\mathcal{E}_{a})$ is spanned by a graph $\mathcal{G}_{b}=(\mathcal{V}_{b},\mathcal{E}_{b})$ when $\mathcal{V}_{a}=\mathcal{V}_{b}$ and $\mathcal{E}_{b}\subseteq\mathcal{E}_{a}$. $\mathcal{G}_{a}$ and $\mathcal{G}_{b}$ can be connected or disconnected.
@@ 2.2 System Graphs and Structure Patterns
3 Modeling of Neuromimetic Dynamic Networks
@@
where $t\in(\tau p,\tau(p+1))$, $p=1,2,\cdots$ and $\tau$ is a time constant denoting the time interval of updating the edge weight. $i,j\in\mathcal{V}$. $c_{n}>0$ represents the decay rate, and $10<c_{a}^{+},c_{a}^{-}<$ 1 exhibits the synaptic depression property. $\phi(\cdot)$ is the activation function. In models of neurodynamics, a commonly used activation function is the sigmoid function. In (1a), $b_{i}=0$ means there is no external input injected to node $i$. The entry $a_{ij}$ represents the connection weight of the pre-synaptic neuron $j$ to the post-synaptic neuron $i$. The function $\gamma_{\theta}$ is designed to emulate neuronal spiking behavior. It is defined as:
@@
where $\theta$ represents the threshold for neuronal activation. It can also be expressed using the standard $ReLU$ function, $\sigma(x)=\max(x,0)$ as $\gamma_{\theta}(x)=-\sigma(-x-\theta)+\sigma(x-\theta)+sign(\sigma(x-\theta))\theta+sign(-\sigma(-x-\theta))\theta$. If $x$ represents the membrane potential of the postsynaptic neuron, the function $\gamma_{\theta}(x)$ models the instantaneous increase in postsynaptic membrane potential caused by the presynaptic release of vesicles, under the condition that the threshold potential $+\theta/-\theta$ is reached. This models the scenario where presynaptic vesicle release results in a discontinuous jump in postsynaptic potential upon crossing the threshold $\theta/-\theta$, which corresponds to the integrate-and-fire model in Burkitt (2006). (1b) models the evolution of connectomes of neurons or brain regions occurring on slower timescales than neuronal state changes, consistent with synaptic plasticity processes that strengthen or weaken synapses over time as studied by Citri and Malenka (2008). Therefore, over a short interval $\tau$, synaptic strengths are fixed and describe the modulation of faster neural state fluctuations described by $x$. Plasticity updates to $A$ occur discretely every $\tau$ units of time. The evolving positive and negative edge weights describe the excitatory and inhibitory plasticity, which correspond to Hebbian and anti-Hebbian rules, respectively. Meanwhile, it is assumed that the synaptic strength is bounded: $\underline{a}^{-}<a_{ij}<\overline{a}^{-}<0$ for all $(j,i)\in\mathcal{E}^{-}$ and $0<\underline{a}^{+}<a_{ij}<\overline{a}^{+}$ for all $(j,i)\in\mathcal{E}^{+}$. For example, when the dynamic of $a_{ij}$ for $(j,i)\in\mathcal{E}^{+}$ evolves to exceed the value $\overline{a}^{+}$, it takes the value $\overline{a}^{+}$; when it is smaller than $\underline{a}^{+}$, it becomes $\underline{a}^{+}$. This assumption is reasonable in neuroscience since synapses are naturally bounded.
@@ 4.1 Boundedness and Stability
The commonly used activation function – sigmoid function satisfies the boundedness assumption. In this paper, we use the range of the sigmoid function to be $[-1,1]$ such that $\sup{\phi}=1$ and $\phi(0)=0$. From the dynamic equations (2), we define $\mathcal{X}=\{\bm{x}\in\mathbb{R}^{n}:|x_{i}|\leq x_{max}\}$, where $x_{max}=\frac{||B\bm{u}||_{\infty}}{c_{n}-d^{in}\max\{\overline{a}^{+},|\underline{a}^{-}|\}}$ denotes the maximum membrane potential.
@@
Since conditions in Lemma 2 establish that the infinity norm qualifies as a Lyapunov function, we omit the detailed proof of the Theorem 3.
@@ 4.3 Resilience and Cascade
It is obtained graphically. From Theorem 4, we obtain that as long as the graph preserves the generalized cactus structure during dynamic evolutions, the system is structurally controllable and observable relative to the same input nodes. Therefore, even if there are neurons and connections that are intermittently available, the system is always structurally controllable and observable. Meanwhile, even if there are edges and nodes unavailable, conditions in lemma 2 still hold. Thus, the state matrix in (2a) is still a diagonally dominant matrix and thus, the system remains stable. ∎